8.28 Ammonia is a colorless gas with a pungent, characteristic odor. It is prepared by the reaction between hydrogen and nitrogen: 3H,(g) N2(g) In a particular reaction, 6.0 moles of NH, were produced. How many moles of H. and how many moles of N, were consumed to produce this amount of NH,?
Answer #1

According to the balance equation 2 moles of NH3 is produced from 3 moles of H2 and 1 mol of N2.

Moles of H2 consumes = (3/2)*6 moles = 9 moles

moles of N2 consumed = (1/2)*6 moles = 3 moles